# Active Advantage-Aligned Online Reinforcement Learning with Offline Data

### v3

@@ Introduction
Fig. 3b compares the performance of $A^{3}$RL to $A^{3}$RL with only advantage in sampling priority (${\sigma}=\exp\left({\mathcal{\xi}\cdot A}\right)$), without density term. The results consistently show that $A^{3}$RL, which incorporates onlineness through the density term $w=d^{\text{on}}/d^{\text{off}}$, outperforms the version that does not. Onlineness measures the likelihood that $A^{3}$RL will experience the given transition during the online exploration and exploitation of the current policy. Transitions experienced during online policy enhancement are more advantageous for policy development. In contrast, focusing on transitions that are unlikely to occur during live interactions with the environment can hinder the progression of policy improvement. This result demonstrates the effectiveness of onlineness term.
Ablation on purely online setting and offline data.
Fig. 3c presents an ablation study comparing regular $A^{3}$RL (in red), purely online $A^{3}$RL (in blue), and SAC (in green), with neither having access to offline data. $A^{3}$RL surpasses its purely online version when utilizing an offline dataset, as the offline data provides a more diverse range of transitions that the online policy might not encounter, effectively demonstrating $A^{3}$RL’s ability to leverage offline datasets. Moreover, the purely online version of $A^{3}$RL outperforms SAC, highlighting $A^{3}$RL’s robustness in environment setting. The results confirm $A^{3}$RL’s effectiveness in a purely online environment and its superiority over SAC in online batch scenarios through active advantage-aligned sampling.
Ablation on priority term.
Fig. 3d presents an ablation study for $A^{3}$RL (in red), where we compare two different sampling strategies: PER as detailed in (Schaul et al. 2015) (named as TD in blue), and a modified version incorporating a density ratio (named as TD+Density in green). The TD-error based sampling strategy prioritizes transitions with larger TD-errors. $A^{3}$RL significantly outperforms both strategies, illustrating that an active advantage-aligned sampling approach is more effective than prioritizing based on TD-error alone. The superior performance of $A^{3}$RL over TD+Density also indicates that prioritizing using the advantage term achieve the better performance compared to the TD-error term.
Conclusion
@@
We present $A^{3}$RL, a novel algorithm for online RL with offline dataset through a confidence-aware active advantage-aligned sampling strategy. This algorithm is theoretically motivated by the objective of shifting the sampling distribution toward more beneficial transitions to maximize policy improvement. We provide theoretical insights for $A^{3}$RL and quantify its enhancement gap. Moreover, we conduct comprehensive experiments with various qualities of offline data, demonstrating that $A^{3}$RL outperforms the SOTA RLPD method with significance. We also conduct multiple ablation studies and confirm the importance of each component within the active advantage-aligned formula and its effectiveness to pure online setting as well. While our approach primarily aims to enhance performance, it may result in higher computational costs due to the calculations needed for determining advantage-aligned sampling priorities. Reducing computational demands will be a focus of our future work.
@@ Appendix A Theoretical Motivation
Proof: (Proof of Theorem 1). Define visitation measures
@@
$=\mathbb{E}_{d^{\widehat{\pi}}}\left[\bigl{\langle}\widehat{\pi},Q^{\widetilde{\pi}}\bigr{\rangle}_{\mathcal{A}}-\alpha\left\langle\widehat{\pi},\log\widehat{\pi}\right\rangle-\bigl{(}\bigl{\langle}\widehat{\pi},\widehat{}Q^{\widetilde{\pi}}\bigr{\rangle}_{\mathcal{A}}-\alpha\left\langle\widehat{\pi},\log\widehat{\pi}\right\rangle\bigr{)}\right$]
@@
See 1 Proof:(Proof of Lemma 1) Under the reparameterization $d^{\text{on}}(a)\propto\exp(\beta_{1}r(a))$ and $\pi^{t+1}(a)\propto\exp(\beta_{2}r(a))$, we have for the coefficient $R^{t}(a;\xi)$ that
@@
Note that the largest probability ratio happens for $\widetilde{a}=\mathop{\rm arg\,max}_{a^{\prime}}r(a^{\prime})$. Since the softmax is strictly less than the argmax when $r$ has different values in each action, the above derivative for action $\widetilde{a}$ is negative, meaning that by increasing $\xi$, the value of $R(\widetilde{a};\xi)$ will decrease. As $\sup_{a}R(a;\xi)=R(\widetilde{a};\xi)$ by our previous discussion, we complete the proof. $\square$
@@ Appendix B Additional Preliminaries
Maximum Entropy RL: Incorporating entropy into the learning objective (as defined in (1)) helps mitigate overconfidence in value estimates, particularly when training with offline datasets. In offline RL, policies may become overly conservative due to limited dataset coverage, leading to suboptimal exploration during fine-tuning. By preserving policy stochasticity, entropy regularization ensures that the agent remains adaptable when transitioning from offline training to online interactions. This controlled exploration has been shown to improve training stability and prevent premature convergence (Ball et al. 2023; Chen et al. 2021; Haarnoja et al. 2018; Hiraoka et al. 2021).
@@ Appendix C Additional Related Work
In an effort to mitigate the sample complexity of online RL (Liu et al. 2023c), offline RL utilizes fixed datasets to train policies without online interaction, however it can be prone to extrapolation errors that lead to overestimation of state-action values. Recent off-policy actor-critic methods (Fujimoto, Meger, and Precup 2019; Kostrikov, Nair, and Levine 2021; Kumar et al. 2020; Wang et al. 2020) seek to mitigate these issues by limiting policy learning to the scope of the dataset, thereby minimizing extrapolation error. Strategies for reducing extrapolation error include value-constrained approaches (Kumar et al. 2020) that aim for conservative value estimates and policy-constrained techniques (Nair et al. 2020) that ensure the policy remains close to the observed behavior in the data. There are several works that leverage advantage estimation to guide policy improvement in purely offline RL, such as LAPO (Chen et al. 2022), A2PR (Liu et al. 2024a), and A2PO (Qing et al. 2024). However, they are not well-suited for online settings because they fail to consider the importance of “onlineness,” measured by the density ratio, to align with the needs of online RL exploration and exploitation. Additionally, they do not account for uncertainty in advantage estimation.
@@
While offline RL methods can outperform the dataset’s behavior policy, they rely entirely on static data (Levine et al. 2020). When the dataset has comprehensive coverage, methods like FQI (Antos, Szepesvári, and Munos 2007) or certainty-equivalence model learning (Rathnam et al. 2023) can efficiently find near-optimal policies. However, in practical scenarios with limited data coverage, policies tend to be suboptimal. One approach to addressing this suboptimality is to follow offline RL with online fine-tuning, however as discussed above, existing methods are prone to catastrophic forgetting and performance drops during fine-tuning (Luo et al. 2023). In contrast, $A^{3}$RL begins with online RL while incorporating offline data to enhance the policy, selectively leveraging offline data to facilitate online policy improvement.
Appendix D Limitations of the prior state-of-the-art.
@@
A drawback of RLPD, as discussed by Ball et al. (2023), lies in its symmetric random sampling method applied to both online and offline data, disregarding the significance of individual transitions for evolving quality of policy. This predefined approach to sampling can potentially lead to less than optimal policy improvements due to the omission of vital data and inefficiencies arising from the use of redundant data. Such inefficiencies fail to offer any positive contribution towards enhancing policy. To address the limitation, our research presents an innovative active data sampling technique, specifically designed to optimize the use of both online and offline data in the process of policy improvement.
Appendix E Experimental Details
@@
Additional experimental results
We explored whether different mixtures of offline datasets can be exploited by $A^{3}$RL. In particular, for the D4RL locomotion: halfcheetah, walker2d and For the Adroit: relocate environments in Fig. 5, mix A corresponds to having the offline dataset consisting of 100% of the -simple Minari dataset, mix B corresponds to 100% -simple and 5% -medium, while mix C corresponds to 100% -simple and 10% -medium. Those proportions were chosen due to the recognizable difference in the performance of RLPD under these different settings. In particular, we observed that all RLPD runs with the offline dataset consisting of 100% of the -simple dataset and no less than 30-40% of the -medium dataset achieve similar performance. Meanwhile at lower percentages such as 5% and 10%, there is a difference between runs of RLPD, which implies that there is significant impact from the offline dataset quality to the bootstrapping from offline transitions.
a halfcheetah mix A
b halfcheetah mix B
c halfcheetah mix C
d walker2d mix A
e walker2d mix B
f walker2d mix C
g relocate mix A
h relocate mix B
i relocate mix C
Figure 5: $A^{3}$RL vs RLPD vs SACfD on (D4RL) halfcheetah, walker2d, and (Adroit) relocate with different offline dataset mixtures. $A^{3}$RL outperforms or performs comparably to RLPD across diverse black-box environments.
For the Adroit environment, relocate, the -cloned dataset plays the role of the -simple dataset above, while the -expert plays the role of the -medium dataset above. The mixtures were generated similarly. $A^{3}$RL robustly outperforms, or at least performs on par with, RLPD across diverse black-box environments.
Additional ablation studies
Ablation on density term.
Fig. 6(a-c) presents further ablation studies on the density term for $A^{3}$RL. We see the distinction in the effectiveness of the density term is more significant over harder tasks like antmaze-medium-play.
a halfcheetah-random (density)
b halfcheetah-medium-replay (density)
c antmaze-medium-play (density)
d halfcheetah-random (online)
e halfcheetah-medium-replay (online)
f antmaze-medium-play (online)
g halfcheetah-random (priority)
h halfcheetah-medium-replay (priority)
i antmaze-medium-play (priority)
Figure 6: Ablation Studies: (a-c): $A^{3}$RL vs $A^{3}$RL without density term. (d-f): $A^{3}$RL vs purely online $A^{3}$RL vs purely online SAC. (g-i): $A^{3}$RL vs PER (TD) vs PER with Density (TD+Density).
Ablation on purely online setting.
Fig. 6(d-f) presents further ablation studies on $A^{3}$RL interacting with the environment in a purely online manner, i.e., the algorithm does not utilize access to offline data. It is consistent throughout tested environments that $A^{3}$RL is able to leverage offline data effectively, especially in harder tasks like antmaze-medium-play where purely online $A^{3}$RL fails to learn in the same number of steps.
Ablation on priority term.
Fig. 6(g-i) presents further ablation studies on the priority term for $A^{3}$RL, where we compare it against the sampling strategy that solely uses TD-error as the priority term, and another that combines the density term with TD-error. The superior performance of $A^{3}$RL over TD+Density over tested environments indicates that prioritizing using the advantage term achieves better performance compared to the canonical TD-error term.
Training and evaluation environments.
@@
Fig. 7 presents snapshots of tested D4RL locomotion tasks: halfcheetah, walker2d, ant and antmaze. halfcheetah, walker2d, and ant have dense rewards, while antmaze has sparse rewards, and all environments are equipped with continuous state and action spaces.
@@
In the halfcheetah environment, the 2D agent resembles a simplified cheetah model with a torso and lined legs, with the objective of forward locomotion and maintaining balance while maximizing speed. In the walker2d environment, the 2D humanoid agent has 2 legs and multiple joints, with the objective of stable bipedal walking without falling. In the ant environment, the agent is a 3D quadrupedal agent with multiple joints and degrees of freedom, with the objective of moving forward efficiently while maintaining balance. For all of these environments, rewards are given for velocity to encourage the agent to move forward efficiently while maintaining balance, and several offline datasets, per (Fu et al. 2020), with varying characteristics, as detailed below, were tested.
@@
Fig. 7(right most) presents snapshots of tested Adroit manipulation tasks: relocate. This environment involve a simulated 28-DoF robotic arm interacting with objects in a 3D space and are characterized by sparse rewards and continuous state and action spaces.
@@
-expert-v2 1M samples from policy trained to completion with SAC
-medium-v2 1M samples from policy trained to 1/3 of expert
-medium-replay-v2 Replay buffer of policy trained to medium
-random-v2 1M samples from randomly initialized policy
Table 1: Locomotion offline dataset.
In the antmaze environment, the aforementioned ant agent is placed in a maze environment and must navigate from a defined start point to a goal. Rewards are binary: $1$ for reaching the goal and 0 otherwise. Varying sizes of the maze were tested: umaze (U-shaped), medium and large; which are naturally also of increasing difficulty.
Figure 7: Environments: halfcheetah, walker2d, ant, antmaze respectively and relocate.
Computing infrastructure and wall-time comparison.
We performed our experiments on a cluster that includes CPU nodes (approximately 280 cores) and GPU nodes (approximately 110 NVIDIA GPUs, ranging from Titan X to A6000, set up mostly in 4- and 8-GPU configurations). On the same cluster, the wall run time of $A^{3}$RL is approximately 1.5 times the run time of regular RLPD and is comparable to Off2On.
Hyperparameters and architectures.

### v4

@@ 6.2 Ablation Studies and Effects of $A^{3}$RL
Ablation on density term. Fig. 3a compares the performance of $A^{3}$RL to $A^{3}$RL with only the advantage term in the sampling priority ($p=\exp\left({\mathcal{\xi}\cdot\underline{A}}\right)$), without the density term, on humanoidstandup-medium. This amounts to setting $\zeta=0$. Our ablation results consistently show that $A^{3}$RL, which incorporates onlineness through the density term $w=d^{\text{on}}/d^{\text{off}}$, outperforms the version that does not. Removing the density term leaves the agent to potentially focus on transitions that are unlikely to occur during online interactions, which can yield sample inefficiencies.
Ablation on advantage term. Fig. 3b compares the performance of $A^{3}$RL to $A^{3}$RL with only the advantage term in the sampling priority ($p=\mathbb{I}^{\text{off}}\underline{{w}}\left({s,a}\right)+\mathbb{I}^{\text{on}}$), without the density term, on hammer-cloned. This amounts to setting $\xi=0$. The results show that $A^{3}$RL with the advantage term surpasses its counterpart that only considers onlineness in prioritizing samples. This superiority is attributed to the advantage term, which effectively screens out transitions that are either non-informative or harmful. For example, even if a transition is near on-policy, it may not necessarily be advantageous for the policy to learn.
Ablation on LCB. Fig. 3c compares the performance of $A^{3}$RL to $A^{3}$RL without the LCB calibration for the advantage term on humanoid-medium. In other words, $\underline{A}=\widehat{A}$, the ensemble mean. This amounts to setting $\beta=0$. This demonstrates the importance of staying pessimistic in advantage estimations, as over-optimism can be compounded by the exponentiation, resulting in a miscalibrated, highly biased sampling scheme.
Pure online setting and TD. Fig. 3d presents an ablation study comparing an online variant of $A^{3}$RL against a pure SAC agent, with neither having access to offline data. The effective prioritization scheme is then $p=\exp(\xi\cdot\underline{A})$, and we still observe the perks of this method, which prioritizes more advantageous transitions. This heuristic is also superior to using the PER TD-error heuristic, as shown in Fig. 3e.
Evolution of offline dataset entropy. The easiest way to quantify the effect of the $A^{3}$RL sampling scheme is to measure the entropy corresponding to the sampling probability on the offline dataset, where the priority has both density and advantage contributions. As shown in Fig. 3f, after the initial phase of RLPD training, this entropy quickly drops off from 13.8 nats, which corresponds to $\approx$ 1M effective transitions, to 10.8 nats, which corresponds to $\approx$ 50K effective transitions. It is via this selection mechanism that $A^{3}$RL is able to achieve better sample efficiency.
Effect of dataset quality and quantity. Even though baseline methods are somewhat comparable to $A^{3}$RL in the high-quality, high-quantity data regime, e.g., Fig. 2a, their performances drastically suffer when either quality or quantity is affected. Fig. 3g demonstrates the robustness of $A^{3}$RL when the dataset is not of expert quality, while Fig. 3h shows that it is still able to learn in a sample-efficient manner with a much smaller offline dataset. In both of these scenarios, all baseline methods are unable to learn, despite both PEX and BOORL having an extensive offline pretraining phase.
7 Conclusion
@@
We present $A^{3}$RL, a novel algorithm for online RL with offline datasets through a confidence-aware active advantage-aligned sampling strategy. This algorithm is theoretically motivated by the objective of shifting the sampling distribution toward more beneficial transitions to maximize policy improvement, and its enhancement gap is quantified. Moreover, we conduct comprehensive experiments with various qualities of offline data, demonstrating that $A^{3}$RL outperforms existing baselines with significance. We also conduct multiple ablation studies and confirm the importance of each component within the active advantage-aligned formula, as well as its superiority in the online setting and over existing prioritization schemes. While our approach primarily aims to enhance performance, it is in essence trading an offline pretraining phase for an extra density network, which could cause higher computational costs with large architectures. Though it is empirically seen that our approach is relatively computationally efficient, further reducing computational demands will be a focus of our future work.
@@ Appendix A Theoretical Motivation
Proof A.1.
(Proof of Theorem 1). Define visitation measures
@@
[…]
@@
See 1
Proof A.2.
(Proof of Lemma 1) Under the reparameterization $d^{\text{on}}(a)\propto\exp(\beta_{1}r(a))$ and $\pi^{t+1}(a)\propto\exp(\beta_{2}r(a))$, we have for the coefficient $R^{t}(a;\xi)$ that
@@
Note that the largest probability ratio happens for $\widetilde{a}=\mathop{\rm arg\,max}_{a^{\prime}}r(a^{\prime})$. Since the softmax is strictly less than the argmax when $r$ has different values in each action, the above derivative for action $\widetilde{a}$ is negative, meaning that by increasing $\xi$, the value of $R(\widetilde{a};\xi)$ will decrease. As $\sup_{a}R(a;\xi)=R(\widetilde{a};\xi)$ by our previous discussion, we complete the proof.
@@ Appendix B Additional Preliminaries
Maximum Entropy RL: Incorporating entropy into the learning objective (as defined in (1)) helps mitigate overconfidence in value estimates, particularly when training with offline datasets. In offline RL, policies may become overly conservative due to limited dataset coverage, leading to suboptimal exploration during fine-tuning. By preserving policy stochasticity, entropy regularization ensures that the agent remains adaptable when transitioning from offline training to online interactions. This controlled exploration has been shown to improve training stability and prevent premature convergence [2, 6, 15, 18].
@@ Appendix C Additional Related Work
In an effort to mitigate the sample complexity of online RL [37], offline RL utilizes fixed datasets to train policies without online interaction, however it can be prone to extrapolation errors that lead to overestimation of state-action values. Recent off-policy actor-critic methods [14, 27, 29, 67] seek to mitigate these issues by limiting policy learning to the scope of the dataset, thereby minimizing extrapolation error. Strategies for reducing extrapolation error include value-constrained approaches [29] that aim for conservative value estimates and policy-constrained techniques [44] that ensure the policy remains close to the observed behavior in the data. There are several works that leverage advantage estimation to guide policy improvement in purely offline RL, such as LAPO [5], A2PR [33], and A2PO [49]. However, they are not well-suited for online settings because they fail to consider the importance of “onlineness,” measured by the density ratio, to align with the needs of online RL exploration and exploitation. Additionally, they do not account for uncertainty in advantage estimation.
@@
While offline RL methods can outperform the dataset’s behavior policy, they rely entirely on static data [31]. When the dataset has comprehensive coverage, methods like FQI [1] or certainty-equivalence model learning [51] can efficiently find near-optimal policies. However, in practical scenarios with limited data coverage, policies tend to be suboptimal. One approach to addressing this suboptimality is to follow offline RL with online fine-tuning, however as discussed above, existing methods are prone to catastrophic forgetting and performance drops during fine-tuning [40]. In contrast, $A^{3}$RL begins with online RL while incorporating offline data to enhance the policy, selectively leveraging offline data to facilitate online policy improvement.
Appendix D Implementation Details of $A^{3}$RL
We provide more extensive details of $A^{3}$RL in Algorithm 1 as Algorithm 2.
Algorithm 2 $A^{3}$RL
4:Initialize offline priority buffer $\mathcal{B}^{\text{off}}$ with offline dataset $\mathcal{D}$, online priority buffer $\mathcal{B}^{\text{on}}=\mathcal{R}\leftarrow\emptyset$ .
5:while True do
6:  Receive initial observation state $s_{0}$
7:  for $t=0,\ldots,T$ do
8:   Take action $a_{t}\sim\pi_{\phi}\left({\cdot|s_{t}}\right)$, update online buffer $\mathcal{B}^{\text{on}}\leftarrow\mathcal{B}^{\text{on}}\cup\left\{\left({s_{t},a_{t},r_{t},s_{t+1}}\right)\right\}$.
9:   for $g=1,\ldots,G$ do
10:     Uniformly sample $\frac{N}{2}$ transitions from $\mathcal{B}^{\text{off}}$ and similarly $\frac{N}{2}$ from $\mathcal{B}^{\text{on}}$. Update $w_{\psi}$ to maximize (5).
11:     Sample with priority $\frac{N}{2}$ transitions from $\mathcal{B}^{\text{off}}$ and similarly $\frac{N}{2}$ from $\mathcal{B}^{on}$ to form batch $b$ of size NN.
12:     Sample set $\mathcal{Z}$ of $Z$ indices from $\left\{1,\ldots,E\right\}$
13:     With $b$, set $y=r+\gamma\Big(\min_{i\in\mathcal{Z}}Q_{\theta^{\prime}_{i}}\left({s^{\prime},a^{\prime}}\right)+\Big$. $\Big.\alpha\log\pi_{\phi}\left({a^{\prime}|s^{\prime}}\right)\Big),a^{\prime}\sim\pi_{\phi}\left({\cdot|s^{\prime}}\right)$
14:     for $i=1,\ldots,E$ do
15:      Calculate importance weights $u_{i}$ via (3).
16:      Update $\theta_{i}$ minimizing loss: $\ell=\sum_{i}u_{i}\cdot\left({y-Q_{\theta_{i}}\left({s,a}\right)}\right)^{2}$
17:     Update target networks: $\theta^{\prime}_{i}\leftarrow\rho\theta^{\prime}_{i}+\left({1-\rho}\right)\theta_{i}$
18:   With $b$, update $\phi$ maximizing objective:
19:     $\frac{1}{E}\sum_{i=1}^{E}Q_{\theta_{i}}\left({s,a}\right)-\alpha\log\pi_{\phi}\left({a|s}\right),$ where $a\sim\pi_{\phi}\left({\cdot|s}\right)$, $\left({s,a}\right)\sim b$.
20:   Update priorities for $b$ according to (6).
Appendix E Limitations of the Prior State-of-The-Art.
@@ Appendix D Implementation Details of $A^{3}$RL
A drawback of RLPD, as discussed by Ball et al. [2], lies in its symmetric random sampling method applied to both online and offline data, disregarding the significance of individual transitions for evolving quality of policy. This predefined approach to sampling can potentially lead to less than optimal policy improvements due to the omission of vital data and inefficiencies arising from the use of redundant data. Such inefficiencies fail to offer any positive contribution towards enhancing policy. To address the limitation, our research presents an innovative active data sampling technique, specifically designed to optimize the use of both online and offline data in the process of policy improvement.
Appendix F Experimental Details
@@
F.1 Additional experiments
a antmaze-umaze
b door-cloned
c door-expert
d door-human
e hammer-human
f pen-cloned
g pen-expert
h pen-human
i relocate-expert
j halfcheetah-medium
k halfcheetah-simple
l hopper-expert
m hopper-medium
n humanoid-expert
o humanoid-medium
p humanoidstandup-medium
Figure 4: Ablation on advantage, $\xi=0$.
a door-cloned
b door-human
c halfcheetah-simple
d hopper-expert
Figure 5: Ablation on lower confidence bound, $\beta=0$.
a door-human
b pen-cloned
c pen-expert
d pen-human
e relocate-expert
f halfcheetah-simple
g hopper-medium
h humanoid-medium
Figure 6: Ablation on online setting.
a antmaze-umaze
b door-expert
c pen-expert
d halfcheetah-simple
e humanoid-expert
f humanoid-medium
Figure 7: Ablation against TD error as priority.
F.2 Environments and datasets.
Environments.
@@ F.1 Additional experiments
Fig. 8 and Fig. 9 present the tested D4RL tasks: the locomotion ones include halfcheetah, walker2d, ant, antmaze, hopper, humanoid; while the Adroit hand ones include pen, relocate, hammer and door. antmaze has sparse rewards, while all other environments have dense reward. All environments are equipped with continuous state and action spaces.
@@
In the halfcheetah environment, the 2D agent resembles a simplified cheetah model with a torso and lined legs, with the objective of forward locomotion and maintaining balance while maximizing speed. In the walker2d environment, the 2D humanoid agent has 2 legs and multiple joints, with the objective of stable bipedal walking without falling. In the ant environment, the agent is a 3D quadrupedal agent with multiple joints and degrees of freedom, with the objective of moving forward efficiently while maintaining balance. humanoid and hopper are similar. For all of these environments, rewards are given for velocity to encourage the agent to move forward efficiently while maintaining balance, and several offline datasets, per [12], with varying characteristics, as detailed below, were tested.
@@
Fig. 8 (right most) presents a snapshot of relocate. This environment involve a simulated 28-DoF robotic arm interacting with objects in a 3D space and are characterized by sparse rewards and continuous state and action spaces. door, hammer, pen are similar high-dimensional robotic arm manipulation tasks.
@@
In the antmaze environment, the aforementioned ant agent is placed in a maze environment and must navigate from a defined start point to a goal. Rewards are binary: $1$ for reaching the goal and 0 otherwise.
Datasets.
We use offline datasets as provided by Minari [69]. For each of the environments above, depending on the type of environment (whether it is under D4RL (Adroit) or mujoco (locomotion)), there are several types of offline datasets. These are detailed in Table 1 and Table 2.
@@
simple 1M steps of agent trained for 1M steps
medium 1M steps of agent trained for 4M steps
expert 1M steps of agent trained for 25M steps
Table 1: Offline dataset descriptions for hopper.
Offline dataset type Description
human 5K steps of human demonstrations
expert 500K steps of fine-tuned RL policy
cloned 500K steps of imitation learning agent using human and expert
Table 2: Offline dataset descriptions for pen.
The exact number of steps may vary per environment, but hopper and pen are representative of the locomotion and Adroit tasks respectively. We refer to Younis et al. [69] for more granular details for each environment.
Figure 8: Environments: halfcheetah, walker2d, ant, antmaze and relocate.
Figure 9: Environments: pen, door, hammer, hopper and humanoid.
F.3 Computing Infrastructure and Wall-time Comparison.
We performed our experiments on a cluster that includes CPU nodes (approximately 280 cores) and GPU nodes (approximately 110 NVIDIA GPUs, ranging from Titan X to A6000, set up mostly in 4- and 8-GPU configurations). On the same cluster, the wall run time of $A^{3}$RL is approximately 1.25 times the runtime of regular RLPD, and is about half the runtime of PEX and BOORL (offline pretraining included). This highlights the superior sample and compute efficiency of $A^{3}$RL.
F.4 Hyperparameters